The position of a point is given as a function of time by x= 4t − 6t + 2t− 1 , where x and t are expressed in meters and seconds. Determine the position, the velocity, and the acceleration of the particle when t= 2 s.

1
Expert's answer
2021-11-16T13:40:39-0500

Given:

"x=4-6t+2t^2"

"t=2\:\rm s"


The position

"x(2)=4-6*2+2*2^2=0"

The velocity

"v(2)=x'(2)=-6+4*2=2\:\rm m/s"

The acceleration

"a(2)=v'(2)=4\:\rm m/s^2"